How much work is required to move a charge of 4nC from a point 2m away to a point 0.5m away from a point charge of 60nC? What is the potential difference between these points?



1
Expert's answer
2021-06-21T11:38:17-0400

(a) We can find the work done required as follows:


"W=kqQ(\dfrac{1}{r_2}-\dfrac{1}{r_1}),""W=9\cdot10^9\ \dfrac{N\cdot m^2}{C^2}\cdot4\cdot10^{-9}\ C\cdot60\cdot10^{-9}\ C\cdot(\dfrac{1}{0.5\ m}-\dfrac{1}{2\ m}),""W=3.24\cdot10^{-6}\ J."

(b) We can find the potential difference as follows:


"W=q\Delta V,""\Delta V=\dfrac{W}{q}=\dfrac{3.24\cdot10^{-6}\ J}{4\cdot10^{-9}\ C}=810\ V."